Comprehensive problem XYZ has the following financial information for 2012: Sales $2M, Net Inc.-3 C.A. $0.4M, F.A. $3.6M C.L. = $0.2 M, LTDE $1 M, C.s-$2M, R.E. = $0.8M What is the sustainable growth rate? 0.4M, Div. $0.1M Numeric Answer:
Answer #1

Sales = $2M

NI = 0.4 M

Div= 0.1 M

C.A = 0.4 M

F.A = 3.6M

C.L=0.2M

LTD =1M

C.S=2M

R.E= 0.8M

ROE = Net Income/ Shareholders’ Equity

= 0.4M/(2M+0.8M) = 0.1429

b) Payout Ratio = Dividends/Netcome

= 0.1M/0.4M = 0.25

Plowback Ratio = 1- Payout ratio

= 1-0.25 = 0.75

Sustainable Growth rate = ROE*Payout ratio/1-ROE*Payout ratio = 0.1429*0.75 / (1 – (0.1429 X0 .75)) = 0.12 = 12.00%